OpenStudy (anonymous):

Please explain how to solve: If x1 and x2 are solutions of the equation 2^4x^2 = 1/4^4x+1 compute the value of x1-x2 Problem 2: Find y when 2^x = 16^y, 16^x = 8^y+5

OpenStudy (anonymous):

Is this the equation for the first part? It's difficult to infer what you mean if it's typed out like that. I'd suggest in the future typing anything complex out using the equation editor in a comment reply. \[2^{4x^2} = \frac{1}{4^{4x}}+1\]

OpenStudy (anonymous):

I'm sorry I didn't know of the equation editor. This is the equation: \[2^{4x ^{2}}=1/(4^{4x+1})\]

OpenStudy (anonymous):

First of all I'd bring the denominator on the right up to the left hand side. Then I'd consider rewriting the 4 (the one that isn't inside an exponent) and see if that helps at all (looks like it will!).

OpenStudy (anonymous):

Thank you for the reply, these are the steps I did. I'm stuck and really have no idea how to solve it. \[2^{4x ^{2}}(4^{4x+1})=1\] \[8^{4x ^{2}+8x+2}=1\] \[\log _{8}(1) = 4x ^{2} + 8x + 2\]

OpenStudy (anonymous):

i think there is a mistake here

OpenStudy (anonymous):

\[4^{4x+1}=(2^2)^{4x+1}=2^{8x+2}\]

OpenStudy (anonymous):

then \[2^{4x^2}\times 2^{8x+2}=2^{4x^2+8x+2}\]

OpenStudy (anonymous):

this tells you \[4x^2+8x+2=0\] or \[2x^2+4x+1=0\] which you can solve using the quadratic formula ,

OpenStudy (anonymous):

Thank you so much, I now see what I was doing wrong.

OpenStudy (anonymous):

oh wait, now i see the problem is something else entirely

OpenStudy (anonymous):

But if I find x won't I be able to just plug in and solve?

OpenStudy (anonymous):

\[2^{4x ^{2}}=\frac{1}{4^{4x+1}}\]

OpenStudy (anonymous):

is that the problem?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

If x1 and x2 are solutions compute the value of x1-x2

OpenStudy (anonymous):

if you write them with the same base of 2, you get \[2^{4x^2}=(2^{-2})^{4x+1}\] \[2^{4x^2}=2^{-8x-2}\]

OpenStudy (anonymous):

now you have to solve \[4x^2=-8x-2\] oh i see, it is the same...

OpenStudy (anonymous):

solve \[2x^2+4x+1=0\]

OpenStudy (anonymous):

The logarithm of 1 in any base is 0. =))

OpenStudy (anonymous):

Thank you both, I got the answer I was looking for.